7 ms·
Is there any (uncontroversial) theory that rigorously defines what a 50 % probability for heads and tails means? It certainly doesn't mean that in the long run
by danbruc 10y ago
Is there any (uncontroversial) theory that rigorously defines what a 50 % probability for heads and tails means? It certainly doesn't mean that in the long run you will obtain the same number of heads and tails because there is a (vanishing) chance that you will always get heads even though the coin is actually fair. And just saying that you will obtain the same or at least similar number of heads and tails with high probability is a circular definition at best.
Does Bayesian thinking just deny the existence of intrinsic probabilities or considers them out of scope? Assuming there is such a thing as true randomness, for example in quantum physics (and it is not due to our ignorance of hidden variables like in Bohemian mechanics), could a Bayesian assign probabilities to the outcomes? Would a Bayesian be mislead into assigning something other than 50/50 for spin up and spin down if he, by chance, only observes spin up although he is really confident that this is just a statistical fluke? If he is mislead, does that mean that uncertainty about the truth of his assumptions creeps in in case the system momentarily deviates from the expected behavior?
- paol 10y ago> doesn't mean that in the long run you will obtain the same number of heads and tails IANAM, but in my understanding that's exactly what it means, if you express "in the long run" as "the ratio tends to 0.5 as the number of tosses tends toward ∞"
- danbruc 10y agoIANAM either, but that seems to be wrong to me. For any finite number of tosses you may certainly not converge to 0.5 even if it becomes very unlikely pretty quickly not to do so. In order for this to be true, it would have to be literally impossible to always get heads on every toss no matter how often you try. But it seems counter-intuitive at the very least that it is inevitable to get tails eventually. What rules out the possibility to get heads over and over again, an infinite number of times? [1] But if it is possible to obtain an infinite number of heads and not a single tails, then it is not true that ratio always converges towards 0.5. In consequence one would have to modify the definition and say that it is only very likely to converge towards 0.5 but that obviously raises the problem that »very likely« sounds a lot like »with high probability« which is the thing we are trying to define, hence this would become a somewhat circular definition. [1] Actually it must not only be impossible to only ever get heads or tails, it must be impossible to obtain any sequence of outcomes that has not exactly the same number of heads and tails. There is only one sequence of outcomes with only heads and only one with only tails but there is already an infinite number of sequences with all heads but one tails or vice versa. And there is an enormous number of possible sequences with 40 % heads and 60 % tails or any other ratio other than 50/50.
- paol 10y ago> For any finite number of tosses you can certainly not converge to 0.5 I think you mean "may not" rather than "can not". But yes, that's why you need to resort to limits[0] to understand the problem. [0] https://en.wikipedia.org/wiki/Limit_of_a_sequence https://en.wikipedia.org/wiki/Limit_of_a_sequence
- danbruc 10y agoThanks, fixed. It understand that it is the limit, but the question I am asking is whether it is possible to always get heads, an infinite number of times. If this is possible, then it is not true that the relative frequency always approaches 0.5 as the number of repetitions goes to infinity. I understand that the probability of obtaining heads an infinite number of times is zero, but does that mean it is impossible? After all any specific infinite sequence of outcomes has probability zero.
- paol 10y ago> the question I am asking is whether it is possible to always get heads, an infinite number of times No. By my definition (lay, may be wrong), the probability is the ratio of heads to tails to which we converge at the limit (infinity). The only way for it to converge to 0 is if the probability of heads was 0 to begin with, which would contradict the initial assertion that it was a fair coin. By the way, from your previous comment it seems to me that you're struggling with getting an intuitive grasp on arguments that rely on infinity. To most people (me included for sure) it's initially difficult to accept that infinity is qualitatively different than a finite amount. Our brain seems to superficially accept the concept, but really it keeps trying to reason about it as a really large (but still finite) quantity. Are you familiar with the argument of whether 0.999... equals 1? (Spoiler: it does) Some people have a really hard time coming to terms with it, and the fundamental problem for many comes down to the same difficulty of reasoning about infinity. Also, it's a fun way to troll people :)
- danbruc 10y ago
- rer0tsaz 10y agoNo, but let me quote Jaynes (Probability Theory: the Logic of Science, 10.10): > ‘When I toss a coin, the probability for heads is one-half.’ [...] the issue is between the following two interpretations: > (A) ‘The available information gives me no reason to expect heads rather than tails, or vice versa – I am completely unable to predict which it will be.’ > (B) ‘If I toss the coin a very large number of times, in the long run heads will occur about half the time – in other words, the frequency of heads will approach 1/2.’ These are not the same except for special circumstances like controlled experiments. Frequentism usually assumes or restricts itself to those special circumstances. The long run here means `for any ε > 0, the probability that the observed frequency n/N lies in the interval (1/2±ε) goes to 1 as N goes to infinity'. There is no such thing as an intrinsic probability, a coin only has a chance of landing heads when tossed. If we know everything about the coin and how it is tossed we could calculate the result. Some ways of tossing a fair coin are biased. Some coins are biased when tossed in fair ways. A fair coin toss exaggerates factors that are difficult to know and control exactly, like the force that we apply on the coin with our finger. Jaynes thinks that `true randomness' such as is postulated by conventional quantum physics is unscientific (10.7). In any case it doesn't matter for calculating, and I don't think many Bayesians lose sleep over whether it's unknown, unknowable or `true randomness'. If the Bayesian has an informative prior, they won't be mislead too much. If a Bayesian is 100% sure beforehand with a δ(x - 0.5) prior they won't be mislead at all, but of course no one is ever 100% sure (Cromwell's rule). On the other hand a frequentist might say p<0.05 and be mislead.
- jordigh 10y ago> Is there any (uncontroversial) theory that rigorously defines what a 50 % probability for heads and tails means? Yeah, Kolmogorov's axioms: https://en.wikipedia.org/wiki/Probability_axioms https://en.wikipedia.org/wiki/Probability_axioms To interpret these axioms for 50% probability means that the measure underneath the density function corresponding to the event "heads" is one-half. But "rigourous" doesn't have anything to do with the natural world. You can't make physics "rigourous", for example, but you can make the mathematics inspired by physics rigourous. Kolmogorov's axioms just give a mathematical description of probability in a formal sense i.e. only discussing its form, not its meaning. Formalism is about saying, "whatever this means, this is how it should behave". It's a very 20th century notion of mathematics.
- danbruc 10y agoI know the Kolmogorov's axioms but I am really more interested in that part they avoid - what is the meaning of a probability of 0.5? It is surly nice that we can operate with probabilities in a (hopefully) self-consistent way, but it bugs me quite a bit that I don't really precisely unterstand what the result of a calculation implies for the real world.
- jordigh 10y agoThe modern approach to mathematics is that there is no "meaning", just like "2" or "derivative" has no meaning. We just say how it behaves, or define it in terms of other things, which eventually bottoms out with undefined terms, such as sets and set membership. This is formalism. How you apply mathematics to the world is not the business of formal mathematics. Whatever you want to do with it is "mere" philosophy. ;-)
- danbruc 10y agoThat is what I meant - I am more asking for a solid philosophical interpretation than a mathematical theory, after all frequentism, Bayesianism and all the other interpretations seem to carry quite a bit of philosophy. On the other hand in another comment it just boiled down to the question whether there is a measure that gives 1 for the set of all infinite binary sequences with 50/50 zeros and ones and 0 for the set of all the other sequences. So it is not pure philosophy what I am interested in.
- kareemsabri 10y ago> Is there any (uncontroversial) theory that rigorously defines what a 50 % probability for heads and tails means? It certainly doesn't mean that in the long run you will obtain the same number of heads and tails because there is a (vanishing) chance that you will always get heads even though the coin is actually fair. It doesn't really require a theory, outside of probability theory itself. According to the law of large numbers, P(x) = 0.5 does imply that, for N trials, the probability of N/2 outcomes being x (heads) approaches 1 as N approaches infinity. And as you pointed out, the probability of all heads is vanishing, as N approaches infinity. P(all heads) = 1 - P(not all heads), and P(all heads) approaches 0 as N approaches infinity, so P(not all heads) must approach 1. The same exercise could be done with the relative distributions of heads and tails, in a large number of trials. Each coin flip is random, but the results will converge to the mean. You will obtain the same number of heads and tails with certainty, not high probability, if you can make N big enough.
- danbruc 10y agoThat means it is totally impossible to get H H H H... or H T T H T T... ad infinitum? I really have no well-founded opinion on that, it just seems very counterintuitive that the coin is not allowed to yield any such sequence, they same as good as any other sequence to me.
- kareemsabri 10y agoNot at all! It yields that sequence all the time, in small numbers. But, it is impossible to get only heads, if you could truly flip a coin infinitely many times, which of course you can't. The larger you make the sequence, the smaller you make the probability of all heads. In your example: P(H,H,H,H) = 0.5 ^ 4 = 0.0625 P(H,H,H,H,H,H,H,H) = 0.5 ^ 8 = 0.00390625 As you can see it's getting pretty small already. However, the probably will never actually reach 0. It's the half step concept, if you keep halving the distance between yourself and an object, you will never actually reach it. But the distance between yourself and the object will approach 0 as the number of steps approaches infinity. 0.5 ^ infinity = 0 Note that P(H|H,H,H) = probability of heads, given 3 previous heads, is still just 0.5. Any independent coin flip is 0.5 chance of being heads. But, when you've done 0 coin flips, and you ask, what are the odds of all heads for the next, say, million flips, they are nearly 0.